Subject: 3.5-m Users Committee minutes 12/08/08

From: Bruce Gillespie

Submitted: Wed, 17 Dec 2008 21:58:13 -0700

Message number: 1121 (previous: 1120, next: 1122 up: Index)

           APO 3.5-m Users Committee Phonecon, 12/08/08

Attending:  Suzanne Hawley, Al Harper, Mark Klaene, Bill Ketzeback,  
Russet McMillan, Jon Fulbright, John Bally, Bruce Gillespie, and  
Michael Strauss

Absent: Scott Anderson, Jon Holtzman, Remy Indebetouw

Minutes taken by Bruce Gillespie


User feedback, comments from institutional representatives:

Princeton (Michael Strauss) - Michael said things are generally fine.   
Princeton TripleSpec users are getting good data while being somewhat  
impacted by some of the telescope problems and new instrument teething  
problems (e.g., rotator, TripleSpec instrument software, discussed  
elsewhere in these minutes).  Some of the troubles with TripleSpec  
were not well backstopped with workarounds or recovery procedures.   
Suzanne said that the issues are currently being actively worked, and  
Mark pointed out that all users should be using the newest version of  
TUI when observing with TripleSpec (Version 1.5.1) because several  
fixes to the instrument-related software have been recently implemented.

Colorado (John Bally) - John reported that the CU users are  
particularly fond of TripleSpec, and said that Mike Skrutskie (the  
TripleSpec instrument PI) will be holding a TripleSpec data-reduction  
seminar in Boulder.

Johns Hopkins (Jon Fulbright) - Jon had nothing to report.

Washington (Suzanne Hawley) - Suzanne noted that there have been  
recent problems with Agile involving instrument software and the  
imminent switchover to TUI control.  She hopes that these issues will  
soon be resolved, noting that Agile is being requested quite a bit in  
the first quarter schedule.

Chicago (Al Harper) - Al mentioned that the UC observers, of which  
there are several regulars, have no complaints.

New Mexico State (Jon Holtzman) - no report

Virginia (Remy Indebetouw) - no report


Discussion of telescope/instruments report - Klaene:

Mark summarized the report, given below, noting the recent spate of  
good weather we've enjoyed.  Some progress has been made with the  
problems with the NA2 rotator, which seem to be largely related to  
electronic communications subtleties.  There have also been a growing  
number of intermittent tertiary rotation failures that seem to be  
caused by a possibly damaged circuit board; we are looking at  
replacing the board, perhaps involving a redesign.


3.5-m Telescope, Instruments, and CIF Projects Highlights, 11/1/08  
through 12/01/08

                               Mark Klaene

0) Overview

Under mostly clear night skies, the telescope has been working well  
during the period.  Issues with the NA2 axis controller, tertiary  
rotation, and TripleSpec caused some loss of observing time.  There  
has been, and continues to be, an active campaign to confront and fix  
these problems--see below for details.  This period also included a  
considerable effort in support of the annual Board of Governors meeting.

1) Telescope

NA2 axis controller: Fritz located a number of problems that have been  
plaguing the NA2 rotator controller.  These include some serial  
communication handshaking and some processing problems.  Recent  
repairs seem to have been very successful.

Tertiary rotation: Nick, with the help of others, found more  
intermittent wiring and also a bad solder joint on a circuit board.   
We are concerned that a major rebuild of the wiring and circuit boards  
for the tertiary rotation system will be necessary before these  
problems are eliminated; we are developing a long term plan for this.

2) Instruments

The DIS-blue ion pump was replaced.  DIS and SPIcam are both fully  
operational with no problems.  DIS focus was checked and found to be  
very good.

Echelle and NIC-FPS both continue to be operational with no problems.   
We are still waiting for UVa to design and build up a new NIC-FPS fan- 
out board that will hopefully address the residual readout noise.

TripleSpec had two different problems during this period.  The first  
had to do with exposure readouts that were all zeros.  Power-cycling  
the array got the instrument working properly.  An investigation is  
underway to determine the cause of these frequent data dropouts.  The  
instrument mounting cart also continues to give us trouble.  A lot of  
minor changes have been incorporated and new mounting "mushrooms" are  
also in work.  We have also started discussions whether the cart needs  
a redesign.  Formal acceptance of the instrument is on hold until  
these problems are fully understood.

Agile is still available at NA2; it will be moved to the TR2 port once  
the new rotator is fully tested.  A problem with the rotator motor  
amplifier has caused a delay of the installation, and the amplifier  
was sent back to the manufacturer.  TUI integration for Agile  
continues.  However, a couple software integration issues have  
developed and are being worked.

A GFP run at the end of the November was completed, and the instrument  
is cold and available.

APOLLO remains operational.  The high-precision gravimeter is now  
slated for arrival later this month, and plans are in progress for its  
installation at the 3.5-m.

3) CIF projects

Funding for next year's CIF projects has been approved by the ARC  
Board of Governors.  The principal projects are direct drives for the  
azimuth and altitude axes, new axis controllers for az and alt, and  
calibration lamp upgrades.  There may be some modest funds available  
to begin studies of new instrumentation and upgrades, depending on  
interest and availability of ARC scientists to lead the effort(s).


Board of Governors meeting - Hawley:

Suzanne reported that the ARC Board of Governors approved the 2008  
Annual Report and the budget request for 2009.  The 2009 budget  
request is reduced relative to the 2008 budget (largely due to a new  
IDC agreement between ARC and NMSU), which enables the direct costs  
for telescope operations and the CIF to be maintained at 2008 levels.   
Suzanne also noted that after the direct-drive az and alt drives  
upgrades are completed, planned capital improvements to the telescope  
itself will be essentially complete.  If the ARC partners can continue  
to support the CIF at current levels, we can begin to address the need  
for new and upgraded instruments, particularly (but not exclusively)  
SPIcam and the echelle spectrograph--we are looking for PIs who can  
propose for and lead new or upgraded instrumentation project(s).   
Additionally, active consideration is being given to adding new  
partners to the consortium, which has been a means in the past of  
acquiring new instruments for the telescope.


TripleSpec block scheduling - Hawley:

This topic was basically a announcement that in light of recent  
teething problems with TripleSpec, we will be using a "partial" block  
scheduling paradigm for TripleSpec in the coming quarter.  This will  
enable expert on-call support for the instrument during the schedule  
blocks.  Users who observe with TripleSpec outside of the schedule  
blocks may have less support available.


NIC-FPS focus - Klaene

It has previously been noted that the NIC-FPS filters, both broad- and  
narrow-band, are not parfocal.  Test data were taken in the past  
month, and were sent to John Bally--John said he will take a look at  
these tests.  Bill Ketzeback said that we need the offset corrections  
to be added to our user documentation, and possibly used by the  
software to make automatic corrections.  Suzanne suggested that this  
be made an action item:

ACTION:  Focus offsets for the various NIC-FPS filter need to be  
determined, documented for the users, and automated corrections  
implemented in the operations software.


Communications while observing, redux - Klaene:

Mark gave an update on progress on developing a VoIP audio  
communication capability between the users and the observatory.  Based  
on recent tests, it appears that Skype may be workable, although more  
experience over extended periods is needed.  Mark asked that we tell  
the users that Skype is now available.  If users are interested, they  
should ask the Observing Specialists for connection details.


Opportunities for new & upgraded instruments - Hawley:

As mentioned above, and last month, Suzanne reiterated that she wants  
to hear if there are people interested in leading new and/or upgraded  
instrument development projects.  So far, Colorado and the University  
of Florida (through our SDSS-III connection) have expressed serious  
interest.  Suzanne added that having new instruments at the 3.5-m  
telescope buys us leverage with other observatories, in the context of  
promoting time-sharing between different facilities.  We hope to have  
CIF funds to partially support instrumentation work.

ACTION:  Anybody with an interest in leading and/or supporting new/ 
upgraded instruments for the telescope should make their intentions  
known to Suzanne Hawley by next month's Users Committee meeting.


Other business:

Suzanne and Russet brought up the issue of target lists given in  
observing proposals, which are not actually used by us for scheduling  
programs on the telescope.  Often these targets are proprietary, and  
users have asked us to edit the target lists out of the proposals  
before they are posted.  Suzanne asked that users not put explicit  
target lists into the proposals; it is too time-consuming for us to  
edit individual proposals.

Bill Ketzeback announced that new TUI catalog files have been added to  
the on-line documentation, and users are encouraged to make use of them.



[open from previous months]:

ACTION: The Users Committee representatives are to ask their  
respective users for suggestions on improving the calibration lamp  
system on the telescope, e.g., flux, wavelength distribution, emission- 
line calibration, etc., and report at the next meeting.

STATUS:  Although more user input would be welcome, this action is  
closed.  We will move forward on improving the blue calibration lamps  
(arcs and continuum) for DIS, beginning in January.

[new actions this meeting]:

ACTION:  Focus offsets for the various NIC-FPS filter need to be  
determined, documented for the users, and automated corrections  
implemented in the operations software.

ACTION:  Anybody with an interest in leading and/or supporting new/ 
upgraded instruments for the telescope should make their intentions  
known to Suzanne Hawley by next month's Users Committee meeting.


Next meeting: The next Users Committee phonecon will be on Monday, 12  
January, at 8:30 AM Pacific Time.  The agenda and other materials will  
be sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1121 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to